The work done in pulling up a block of weighing 2KN for a length of 10m on a smooth plane inclined at angle of 15 with the horizontal is
1
Expert's answer
2011-10-07T11:38:42-0400
Block is rising up to the height H = 10m * sin15 = 2.588m.
So, the work will be W = F*H = 2000N * 2.588m = 5176J.